A ball is thrown into the air from the balcony of an apartment building and falls to the ground. The height, h, in metres, of the ball relative to the ground, t seconds after being thrown is given by: h=-5t^2 + 10t + 35
c)After how many seconds does the ball hit the ground?
*I know that the equation in vertex form is: h=-5(t-1)^2+40

Answer by stanbon(75887)

A ball is thrown into the air from the balcony of an apartment building and falls to the ground. The height, h, in metres, of the ball relative to the ground, t seconds after being thrown is given by: h=-5t^2 + 10t + 35
*I know that the equation in vertex form is: h=-5(t-1)^2+40
-----
So the vertex (top point) is (1,40)
-----
c)After how many seconds does the ball hit the ground?
Solve::
-5(t-1)^2 + 40 = 0
-5(t-1)^2 = -40
(t-1)^2 = 8
----
t-1 = 2sqrt(2)
----
Ans: t = [1+2sqrt(2)] seconds
----
Cheers,
Stan H.
---------------